3. A pendulum is released from an initial height 83.0 cm (a) above its lowest point. It then swings down to its lowest point (b), where it contacts a horizontal rod that interrupts its motion and causes it to swing upward with a new radius, as shown below. The highest point it can swing to is 78.2 cm (c). What will be its speed at tha point?
Answer #1

Here ,

for the speed at the lowest point

before the collision ,

u = sqrt(2 * g * h)

u = sqrt(2 * 9.8 * 0.83)

u = 4.03 m/s

after the collision

speed , v = sqrt(2 g * h2)

v = sqrt(2 * 9.8 * 0.782)

v = 3.915 m/s

the velocity at that point after collision is 3.915 m/s